Shift report 20010919 16:00-24:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Wed Sep 19 2001 - 00:10:36 EDT

  • Next message: Apache: "Shift report 20010919 00:00-08:00"

    ------------------------------------------------------------
    Date: Wed Sep 19 00:10:36 2001, Shift: 16:00-24:00
    Supervisor: Y.K.Lee, with: 
    Log book#9 pages 4 - 7
    
    Shift Summary:
    ------------------------------------------------------------
    1.After various problems and ramp failure, they tried to
      fill with 6 bunches as a test ramp but blue quench
      ensued. No beam through the shift.
    2.FV modified the fiber map for scintilating calibration
      fibers in MRS.
    3.D5 display on pet mulfunctioned. Magnet technicians cannot
      fix it. They left the magnet at 700A by measuring the
      voltage drop across the shunt and determining the current
      to be steady.
    4.Magnets are now set and ready to run
        FS at 8 degrees, 1/5 A
        MRS at 90 degrees, 4B
    
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b30 : Wed Sep 19 2001 - 00:11:22 EDT